OpenStudy (anonymous):

How would I solve the following trig question 4cosx+3sinx into the form c sin(x-a)

OpenStudy (anonymous):

Let: \[r \cdot \sin(a) = 4\] \[r \cdot \cos(a) = 3\] \[\implies r \cdot \sin(a) .\cdot \cos(x) + r \cdot \cos(a) \cdot \sin(x) \implies r(\sin(a) \cos(x) + \cos(a) \sin(x))\] \[\implies \color{green}{r \cdot \sin(x + a)}\]

OpenStudy (anonymous):

We require negative in middle..

OpenStudy (anonymous):

I think c is 5

OpenStudy (anonymous):

or r

OpenStudy (anonymous):

r is 5 here.. \[r^2 = \sqrt{4^2 + 3^2} \implies r = 5\]

OpenStudy (anonymous):

The issue I am having is finding A

OpenStudy (anonymous):

Divide both..

OpenStudy (anonymous):

I know in my book is say a is 5.3559 but I am not sure how.

OpenStudy (anonymous):

\[\frac{r \cdot \sin(a)}{r \cdot \cos(a)} = = \tan(a) = \frac{4}{3} \implies a = \tan^{-1}(\frac{4}{3})\]

OpenStudy (anonymous):

I get .9273 something must be wrong.

OpenStudy (anonymous):

I got 53.13..

OpenStudy (anonymous):

Yeah I think because my calculator is in radians.

OpenStudy (anonymous):

in degrees I got 53.13 too.

OpenStudy (anonymous):

It is in radians and not degrees..

OpenStudy (anonymous):

Yes the answer is asked for in radians.

OpenStudy (anonymous):

So this is the answer in my opinion if I have done right.. And this is in radians..

OpenStudy (anonymous):

Maybe my book has a typo or something because your answer makes sense.

OpenStudy (anonymous):

Yes there may be a misprint..

OpenStudy (anonymous):

well thanks anyway.

OpenStudy (anonymous):

Welcome..

OpenStudy (phi):

I would start by noting sin(x-a)= sin(x) cos(a) - cos(x) sin(a) and c sin(x-a) is c cos(a) sin(x) + -c sin(a) cos(x) match the coefficients to 3 sin(x) + 4 cos(x) we get c cos(a)= 3 and -c sin(a)= 4 or tan(a)= -4/3 the inverse tan of -4/3 is -53.13 degrees or rather 306.87 degrees or 5.356 rads also, we get c from (c cos(a))^2 + ( -c sin(a))^2= 9+16= 25 c^2 ( cos^2(a) + sin^2(a))= 25 c= 5 we get 3 sin(x) + 4 cos(x)= 5 sin(x-5.356)

OpenStudy (anonymous):

I am wrong.. @phi

OpenStudy (phi):

no, you said "We require negative in middle.." you just got side tracked...